Acetic acid has a Ka = 1.78×10-5. Determine the pH of a 0.250 M solution of
acetic acid
1
Expert's answer
2020-05-12T11:35:02-0400

The balanced equation of the reaction of the dissociation of acetic acid is:

HCOOH + H2O "\Leftrightarrow" H3O+ + HCOO-.

The pH is defined as:

"pH = -log[H_3O^+]" .

The initial concentration of HCOOH is 0.250 M. After the establishment of the equilibrium, the concentration of HCOOH is 0.250 -x, and the concentrations of H3O+ and HCOO- are equal to x. Then, the equilibrium constant is expressed as:

"K_a = \frac{x·x}{0.250-x}" .

The equilibrium concentration of H3O+ is:

"x^2 + K_a\u00b7x - 0.250\u00b7K_a =0"

"x^2 + 1.78\u00b710^{-5}\u00b7x - 0.250\u00b71.78\u00b710^{-5} =0"

"x = 0.00210" M.

"[H_3O^+] = 0.00210" M.

The pH of the solution is:

"pH = -log(0.00210) = 2.68" .

Answer: the pH of a 0.250 M solution of acetic acid is 2.68.